OpenStudy (anonymous):

What is 5/16 written as a percent?

OpenStudy (anonymous):

5/16 * 100 = 500/16 = 31.25%

OpenStudy (anonymous):

Thank you

OpenStudy (igreen):

@NandagopalPV Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (anonymous):

Oh, sorry. I didn't know that.